OpenStudy (anonymous):

suppose you have 100 ft of string to rope off a rectangular section for a bake sale at a school fair. The function A= -x^2+50x gives the area of the section in square feet, where x is the width in feet. what width gives you the maximum area you can rope off? What is the maxium area? what is the range of the function?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

make a square

OpenStudy (anonymous):

is that all @misty1212

OpenStudy (misty1212):

yes you want the math teacher explanation?

OpenStudy (anonymous):

yes please @misty1212

OpenStudy (misty1212):

the vertex of \(A= -x^2+50x\) is found by computing \(-\frac{b}{2a}\) which in your case it \[-\frac{50}{2\times (-1)}=25\]

OpenStudy (misty1212):

you have 100 feet of fence, each side should be 25 to maximize the area, i.e. it should be a square this is more or less obvious, the rectangle with fixed perimeter has the largest area if you make it a square a square is the most symmetric rectangle

OpenStudy (anonymous):

thank you @misty1212

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]